OpenStudy (anonymous):

What is the slope of the line? y = -5x -9

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@shortycme

OpenStudy (anonymous):

@J-bird

OpenStudy (zale101):

-5 is the slope

OpenStudy (anonymous):

the slope i think is -5

OpenStudy (zale101):

member, the slope intercept form y=mx+b?

OpenStudy (anonymous):

ok thx /) /) ( ^_^) C (") (")

JoelTheBoss (joel_the_boss):

Slope= " y=mx+b"

OpenStudy (nincompoop):

no slope is not y=mx+b slope is the m in y = mx+b

OpenStudy (anonymous):

your correct @nincompoop

OpenStudy (anonymous):

@nincompoop is that you in the picture? if so what dress is that?

OpenStudy (nincompoop):

slope, \( \sf \huge m = \frac{\Delta y}{\Delta x} = \frac{y_2-y_1}{x_2-x_1}\) is the measure of steepness of a line

OpenStudy (anonymous):

ok thx nincompoop